a rectangular vacant lot is 100 feet long. it's area (in square feet) is twice its perimeter (in feet). how wide is the lot?

Answer by nerdybill(7384)

a rectangular vacant lot is 100 feet long. it's area (in square feet) is twice its perimeter (in feet). how wide is the lot?
Let w = width
then
area is 100w
perimeter is 2(100+w) = 200+2w
.
From:"it's area (in square feet) is twice its perimeter" we get our equation:
100w = 2(200+2w)
dividing both sides by 2:
50w = 200+2w
48w = 200
w = 200/48
w = 4.167 feet